OpenStudy (anonymous):

The function b(x) = 2x − 5 determines how many bags of dog food need to be purchased for an animal shelter, where x is the number of dogs at the shelter. The shelter manager uses m(b(x)) to find the amount of money to bring for the dog food purchase. The function m(x) = 3x + 1. Solve for how much money to bring when there are 10 dogs in the shelter. 15 31 46 52

ganeshie8 (ganeshie8):

write down the given functions

ganeshie8 (ganeshie8):

# of bags required :- b(x) = 2x − 5 amount of money required for x bags:- m(x) = 3x + 1

ganeshie8 (ganeshie8):

amount of money required for x dogs :- m(b(x))

OpenStudy (anonymous):

so we would insert 10 into the equation in place of x because there is 10 dogs?

ganeshie8 (ganeshie8):

yup, you have 3 functions, whuch function gives u money, if u insert # of dogs ?

OpenStudy (anonymous):

last 1

OpenStudy (anonymous):

m(x) = 3x + 1

ganeshie8 (ganeshie8):

m(x) = 3x + 1 gives u money, if u enter # OF BAGS

OpenStudy (anonymous):

so we need to use m(b(x)) = 3x + 1?

ganeshie8 (ganeshie8):

but we dont ahve # of bags, we just have # of dogs

ganeshie8 (ganeshie8):

we need to use :- m(b(x))

ganeshie8 (ganeshie8):

put x = 10 m(b(10))

ganeshie8 (ganeshie8):

find b(10) first

ganeshie8 (ganeshie8):

b(x) = 2x − 5 b(10) = 2(10) − 5 = 15 put this in m(b(10))

ganeshie8 (ganeshie8):

m(b(10)) m(15)

ganeshie8 (ganeshie8):

find m(15) now

ganeshie8 (ganeshie8):

m(x) = 3x + 1 m(15) = ?

ganeshie8 (ganeshie8):

u wid me still ? :)

OpenStudy (anonymous):

yeah I gotcha

OpenStudy (anonymous):

m(15) = 3(15) + 1 45+1 m(15) = 46

ganeshie8 (ganeshie8):

good job !

OpenStudy (anonymous):

so they need 46 dollars for dog food :)

ganeshie8 (ganeshie8):

yup ! u like dogs... ur dp says it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

my picture is my dog :)